OpenStudy (anonymous):

f(x) = 3(x – 2)^2(x + 5) find the zeroes

OpenStudy (mathmate):

Hints: The graph of the function crosses the x-axes at x=-5 (from x+5=0) and is tangent to the x-axis at x=2 (from (x-2)^2=0). Can you find the zeroes?

OpenStudy (anonymous):

help me determine the zeroes

OpenStudy (mathmate):

Hint: A zero of a polynomial is where the curve crosses the x-axis or where it is tangent to the x-axis.

OpenStudy (anonymous):

th6anks i understand now

OpenStudy (anonymous):

would the y-intercept be 60?

OpenStudy (mathmate):

yes, exactly. the polynomial expands to 3x^3+3x^2-48x+60, and the constant term =60 is the y-intercept.